# EEG-ImageNet: An Electroencephalogram Dataset and Benchmarks with Image Visual Stimuli of Multi-Granularity Labels

### v1

@@ A.3 Experimental Setup Details
In the object classification task, we conduct experiments under three different granularity settings: the "all" task includes all 80 categories; the "coarse" task includes 40 coarse-grained categories; and the "fine" task includes 8 fine-grained categories that belong to the same parent node, with the average accuracy calculated across 5 groups. Each classification model maintain parameter consistency across the three tasks. We train one model per participant, ensuring that the parameters are consistent between models for different participants as well.
The model structures and hyperparameters are as follows. For SVM, we try linear, polynomial, and radial basis function (RBF) kernels. The regularization parameter is tested from values {}. For RandomForest, we try to set the number of trees in the forest from values {}, with all other parameters set to their default values. For KNN, we set the number of neighbors to . For ridge regression, all parameters are set to their default values. For RGNN, when calculating the edge weights between electrodes, we use the hardware parameters of our data collection device to determine the topological coordinates of each electrode. In addition to the standard implementation, we add two batch normalization layers. The main hyperparameters adjusted are the number of output channels of the graph convolutional network (i.e., the hidden layer dimension) and the number of hops (i.e., the number of layers). These are set to and respectively. For EEGNet, we use the standard implementation and set the length of the first step convolution kernel to half the number of sampling time points, which is 200. The main hyperparameters adjusted are the number of output channels for the first convolutional layer (F1) and the depth multiplier (D), which are set to and respectively. For MLP, we set two hidden layers with dimensions of 256 and 128, respectively. Each linear layer is followed by a batch normalization layer and a dropout layer with a probability of 0.5.
@@
In the image reconstruction task, we first use blip-image-captioning-base to obtain captions for each stimulus image, with parameters set to "max length": 200 and "num beams": 20 to achieve relatively detailed descriptions. Then, we use the CLIP ViT-L/14 version’s tokenizer and text encoder to compute the CLIP embedding for each caption, resulting in dimensions of (77, 768). Next, we map the frequency domain features of the EEG signals to (77, 768) using an MLP with two hidden layers, having dimensions of 1024 and 2048, respectively. This output is then fed into a frozen Stable Diffusion 1.4 model built with the diffusers library. After 50 steps of inference, we obtain image outputs with dimensions of 512x512. The Stable Diffusion model we built specifically consists of the UNet2DConditionModel and AutoencoderKL, and we use PNDMScheduler as the noise scheduler.
Two-way identification refers to a method where the generated image is compared with the original image and one distractor image. The task is to correctly identify the original image from the pair. This method evaluates the ability of the reconstruction pipeline to produce images that are distinguishable and recognizable as the original stimuli. In this study, we selected the second and fifth convolutional layers of AlexNet, the output before the linear layer of Inception, and the embeddings from CLIP ViT-L/14 as comparison features to calculate two-way identification.
@@ A.4 Additional Experimental Results
Table 5 shows the performance of the best-performing participant across all models and tasks.
@@
Figure 4 shows the accuracy for each participant in the object classification task across SVM, MLP, and RGNN models. We find that the ranking of participants’ accuracy is relatively consistent across different models.
Figure 5 presents more image generation results selected from other participants, with Figure 5(a) showing good cases and Figure 5(b) showing bad cases. We identified three main types of bad cases. Similar to the first two images, the reconstructed images lack or misrepresent low-level information such as color and shape. These errors are relatively common and are due to the limitations of our feature mapper and the simple structure of the reconstruction pipeline, resulting in insufficient information restoration. Similar to the latter two images, the reconstructed images lack detail. This limitation is due to the number of denoising steps in the diffusion model and the inherently low signal-to-noise ratio of EEG signals.
We also observed that for certain categories, especially fine-grained ones, all test data points resulted in near-noise outputs, which drew our attention. When we directly input category labels as text prompts into Stable Diffusion 1.4, we found that the generated images had poor realism and three-dimensional structure. Figure 5(c) compares these images with those generated by our reconstruction pipeline from the training set. This improvement suggests that we can use EEG, which can be quickly and extensively obtained as human feedback signals, to enhance the performance of text-image pre-trained models or generative models. This will be the direction of future research.
@@ A.5 Temporal Effect
In our experimental paradigm, to reduce the cognitive load on participants, we group images of the same category together and use the RSVP (Rapid Serial Visual Presentation) paradigm for continuous rapid stimulation. This approach may cause the model to learn temporal continuity features rather than the intrinsic characteristics of the category stimuli. In this study, we use narrower cropped time segments as input features and divide the training and test sets based on temporal order to reduce the impact of temporal effects. In Figure 6, we plotted the average classification accuracy for images at different index positions in the test set under various training and test set splits.
We observed that the first few images in the test set have significantly higher accuracy, indicating a strong temporal effect, while the accuracy tends to stabilize for the subsequent images. Next, we also used several statistical methods for analysis. For the 30-20 split, we first calculated the sliding window standard deviation to measure local volatility. With a sliding window size of 5, the standard deviation at the 14th data point was less than 0.01, indicating a convergence trend at this point. We then checked the stationarity of the data using the ADF test, obtaining an ADF statistic of -1.1874 and a p-value of 0.6790, suggesting that the data might not have fully converged. For the 20-30 split, the standard deviation at the 16th data point was less than 0.01 with a sliding window size of 5. The ADF test yielded a statistic of -3.8505 and a p-value of 0.0024, indicating that the data is stationary and convergent.
In summary, EEG-ImageNet is subject to certain temporal effects. The methods we employed reduced this impact but did not completely eliminate it. Therefore, we recommend that all researchers using this dataset adopt preprocessing techniques similar to ours.

### v2

@@ A.3 Experimental Setup Details
We conduct experiments under three different granularity settings: the ”all” task includes all 80 categories; the ”coarse” task includes 40 coarse-grained categories; and the ”fine” task includes 8 fine-grained categories that belong to the same parent node, with the average accuracy calculated across 5 groups.
The model structures and hyperparameters are as follows. For SVM, we try linear, polynomial, and radial basis function (RBF) kernels. The regularization parameter is tested from values {}. For RandomForest, we try to set the number of trees in the forest from values {}, with all other parameters set to their default values. For KNN, we set the number of neighbors to . For ridge regression, all parameters are set to their default values. For RGNN, when calculating the edge weights between electrodes, we use the hardware parameters of our data collection device to determine the topological coordinates of each electrode. In addition to the standard implementation, we add two batch normalization layers. The main hyperparameters adjusted are the number of output channels of the graph convolutional network (i.e., the hidden layer dimension) and the number of hops (i.e., the number of layers). These are set to and respectively. For EEGNet, we use the standard implementation and set the length of the first step convolution kernel to half the number of sampling time points, which is 200. The main hyperparameters adjusted are the number of output channels for the first convolutional layer (F1) and the depth multiplier (D), which are set to and respectively. For MLP, we set two hidden layers with dimensions of 256 and 128, respectively. Each linear layer is followed by a batch normalization layer and a dropout layer with a probability of 0.5. In the PT task, during the pre-training phase we train with half of the learning rate for half of the epochs.
@@ A.4 Additional Experimental Results
Table 5 shows the average results of all participants in Stage 1 and Table 6 shows the average results of all participants in Stage 2.
| Model | Acc (all) | Acc (coarse) | Acc (fine) | F1 (all) | F1 (coarse) | F1 (fine) | |
| Classic model | Ridge | 0.286±0.074† | 0.394±0.081† | 0.583±0.074† | 0.261±0.070† | 0.373±0.082† | 0.610±0.121† |
| KNN | 0.304±0.086† | 0.401±0.097† | 0.696±0.068† | 0.286±0.081† | 0.380±0.096† | 0.717±0.132† | |
| RandomForest | 0.349±0.087† | 0.454±0.105† | 0.729±0.072† | 0.323±0.083† | 0.425±0.099† | 0.723±0.092† | |
| SVM | 0.392±0.086† | 0.506±0.099† | 0.778±0.054† | 0.378±0.083† | 0.486±0.105† | 0.770±0.054† | |
| Deep model | MLP | 0.404±0.103† | 0.534±0.115 | 0.816±0.054 | 0.397±0.100† | 0.523±0.108 | 0.819±0.053 |
| EEGNet* | 0.260±0.098† | 0.303±0.108† | 0.365±0.095† | 0.251±0.095† | 0.291±0.098† | 0.374±0.102† | |
| RGNN | 0.405±0.095 | 0.470±0.092† | 0.706±0.073† | 0.401±0.098 | 0.455±0.087† | 0.723±0.079† | |
| Model | Acc (all) | Acc (coarse) | Acc (fine) | F1 (all) | F1 (coarse) | F1 (fine) | |
| Classic model | Ridge | 0.182±0.053† | 0.253±0.074† | 0.431±0.108† | 0.178±0.052† | 0.243±0.075† | 0.438±0.107† |
| KNN | 0.220±0.081† | 0.310±0.113† | 0.574±0.119† | 0.211±0.083† | 0.299±0.105† | 0.565±0.134† | |
| RandomForest | 0.268±0.101† | 0.358±0.129† | 0.609±0.136† | 0.259±0.098† | 0.341±0.117† | 0.596±0.139† | |
| SVM | 0.281±0.090† | 0.368±0.107† | 0.657±0.140† | 0.271±0.084† | 0.365±0.109† | 0.648±0.134† | |
| Deep model | MLP | 0.297±0.093 | 0.395±0.110 | 0.718±0.149 | 0.285±0.087 | 0.392±0.108 | 0.710±0.140 |
| EEGNet* | 0.169±0.044† | 0.244±0.095† | 0.377±0.096† | 0.160±0.041† | 0.228±0.088† | 0.372±0.096† | |
| RGNN | 0.302±0.097 | 0.401±0.105 | 0.693±0.140† | 0.297±0.100 | 0.388±0.106 | 0.701±0.142† | |
Table 7 shows the performance of the best-performing participant across all models and tasks of Stage 1..
@@
Figure 5 shows the accuracy for each participant in the object classification task of Stage 1 across SVM, MLP, and RGNN models. We find that the ranking of participants’ accuracy is relatively consistent across different models.
@@ A.5 Temporal Effect
In Figure 6, we plotted the average classification accuracy for images at different index positions in the test set under various training and test set splits to show the temporal effect in Stage 1.
We observed that the first few images in the test set have significantly higher accuracy, indicating a strong temporal effect.
A.6 The Use of Large Language Models
In this work, we leveraged large language models (LLMs) to assist in manuscript preparation, including refining the text for clarity and style, as well as facilitating literature retrieval. All LLM-generated suggestions were carefully reviewed, edited, and integrated by the authors to ensure scientific accuracy and consistency with our own writing voice. We acknowledge the ongoing discourse around the ethical use of LLMs in scholarly writing—particularly regarding transparency, originality, and accountability. We transparently report the use of LLM assistance and reaffirm that all substantive intellectual contributions (e.g. experimental design, data analysis, interpretation) originated from the authors.